OpenStudy (anonymous):

A housekeeping service charges $20 plus an hourly rate. The service costs $50 for 3 hours of work and $70 for 5 hours of work.

undeadknight26 (undeadknight26):

What is the question?

OpenStudy (anonymous):

a) Write an equation to show how the cost (c) is a linear relationship to the number of hours of work (h). b) How much does 13 hours of work cost?

OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!

OpenStudy (lena772):

(a) So she makes $50 for 3 hours. The fixed rate charged is $20. So automatically we know to subtract that from the total charge. 50-20=30. So we want to find the rate paid per hour besides the fixed rate. They charge $30 for 3 hours without the fixed rate so for the rate per hour we would divide 30 by 3. This gives us $10 per hour. c(h)=20+10h We are going up by 10 every time, therefore it is linear.

OpenStudy (lena772):

b) Just substitute 13 into the equation where h is. c(13)=20+10*13 c(13)=20+130 c(13)=150

OpenStudy (anonymous):

absolutely right @Lena772 :) great work !

OpenStudy (lena772):

Thank you @annas